OpenStudy (anonymous):

the points (x, 3) and (-2, 2) have a slope of 1/4. What is the x-coordinate of the point (x, 3)?

ganeshie8 (ganeshie8):

slope = (y2-y1)/(x2-x1) 1/4 = (2-3)/(-2-x) now can you solve for x ?